### v2

# Benchmarking LLM-based Machine Translation on Cultural Awareness

Abstract
Translating cultural-specific content is crucial for effective cross-cultural communication. However, many MT systems still struggle to translate sentences containing cultural-specific entities accurately and understandably. Recent advancements in in-context learning utilize lightweight prompts to guide large language models (LLMs) in machine translation tasks. Nevertheless, the effectiveness of this approach in enhancing machine translation with cultural awareness remains uncertain. To address this gap, we introduce a new data curation pipeline to construct a culturally relevant parallel corpus, enriched with annotations of cultural-specific items [^1] The code and data will be released at: https://github.com/BigBinnie/Benchmarking-LLM-based-Machine-Translation-on-Cultural-Awareness. Furthermore, we devise a novel evaluation metric to assess the understandability of translations in a reference-free manner by GPT-4. We evaluate a variety of neural machine translation (NMT) and LLM-based MT systems using our dataset. Additionally, we propose several prompting strategies for LLMs to incorporate external and internal cultural knowledge into the translation process. Our results demonstrate that eliciting explanations can significantly enhance the understandability of cultural-specific entities, especially those without well-known translations.
1 Introduction
Despite the notable success of MT systems in recent years, thanks to multilingual pre-trained language models Aharoni et al. (2019), their performance when translating cultural-specific items (CSIs) remains poor. This is primarily due to the gap between the cultural differences associated with languages Akinade et al. (2023); Liebling et al. (2022). The variation of cultural-specific items across cultures Woolford (1983) has been a long-standing aspect of translation studies Newmark (1988); Persson (2015); Fernández Guerra (2012). However, existing terminology translation datasets and methods mostly focus on popular domains (e.g., medicine, finance) Dinu et al. (2019); Ghazvininejad et al. (2023), yet do not cover much on cultural-specific aspects. Distinguished from general terms, cultural-specific items are unique to particular cultural groups, making their literal translations difficult for people from other cultures to understand. Moreover, many CSIs have no translations into other languages, further increasing the difficulty of data collection and evaluation of the translation performance. For example, in Figure 1, a Taiwanese dish called “五更肠旺”, which is chitterling in hot pot, has no well-known existing English translations. The translation of ChatGPT “Wu Geng Intestine Soup” is misleading and can not be easily understood by English native speakers. In contrast, a professional human translator would elicit an explanation to make the translation more understandable for readers. As a result, the nuanced characteristics of CSIs and the limited availability of data sources pose challenges in building culture-aware MT systems that can translate these terms into understandable ones in target languages.
Figure 1: Cultural translation errors made by Google Translate and ChatGPT systems
Recently, a new translation paradigm has emerged, which employs prompts to guide large language models (LLMs) to perform machine translation Brown et al. (2020) in a zero-shot or few-shot fashion. With this flexible paradigm, cultural knowledge can be seamlessly incorporated into LLM translation prompts. However, LLM-based translation is sensitive to prompting strategies and prone to generating hallucinations Ji et al. (2023). As shown in Figure 1, the translation of “A Few Good Men” as “壮志凌云” (i.e., Top Gun) wrongly refers to a completely different film. Given that, the cultural sensitivity of LLM-based MT systems remains an open question, especially when dealing with culturally relevant content. Moreover, there is scant research comparing LLM-based translation versus NMT systems regarding their cultural awareness, partly due to the absence of a rich-annotated culturally sensitive parallel corpus and reliable automated metrics for evaluation. This scarcity further hinders the MT development in navigating cultural nuances, limiting the MT systems’ efficacy in promoting cross-cultural communication.
In this study, we focus on evaluating the culture-awareness of MT systems, addressing two research questions: (1) NMT v.s. LLM-based MT, which group of MT systems embed more cultural awareness? (2) Diving into LLM-based MT, what prompting strategy benefits the most for guiding LLMs in cultural-specific translation? To this end, we propose a novel data curation pipeline to build a culture-centered parallel corpus with limited human annotation efforts. To capture challenging examples with geo-metadata, we further conduct cultural knowledge augmentation for CSIs. Our parallel corpora include 6 language pairs (i.e., English-Chinese, English-French, English-Spanish, English-Hindi, English-Tamil, and English-Telugu), covering 7,253 CSIs in 18 concept categories from over 140 countries and regions.
To assess the cultural nuance, we design new automatic metrics targeting the translation quality of cultural concepts, both on the translation accuracy and understandability. Different from existing evaluation metrics which mainly focus on semantic similarity with the reference Anastasopoulos et al. (2021), we define a new metric called understandability to evaluate the CSI translation in a reference-free manner. We use our metrics to compare a series of MT systems, including LLM-based MT, traditional NMTs, and commercial MT systems, on translating cultural content. Moreover, we examine a set of prompting strategies to endow cultural knowledge into LLM-based MT. The traditional terminology translation method enhances the accuracy of translating CSIs that have well-known references but falls short of translating no-translation CSIs. However, incorporating CSI explanations in the prompt significantly improves the translation quality, especially for no-translation CSIs. Our further investigation into different MT systems’ translation strategies reveals that employing more sophisticated approaches, such as describing CSIs and substituting them with similar content in the target language, effectively improves the readability of CSI translations.
2 Culturally Relevant Data Construction
Figure 2: A data curation pipeline for constructing a cultural-specific English-Chinese parallel corpus.
A wide range of well-structured multilingual knowledge gathered by various crowd workers in Wikipedia makes it an ideal source for collecting CSIs and their descriptions in aligned languages. The overall workflow of our data collection includes four stages: (1) building a wiki-centered cultural taxonomy (§2.1); (2) curating parallel entities related to diverse sociocultures (§2.2); (3) augmenting geo-metadata (§2.3). The curation pipeline with one example is shown in Figure 2. We have more data examples in Appendix A.
2.1 Cultural Taxonomy Extraction
Given that culture is an abstract concept, it is hard to directly capture fine-level cultural characteristics from texts. With this consideration in mind, we referred to an existing CSI classification framework  Newmark (1988), which has been popularly used in the study of human translations of cultural concepts, to identify culturally relevant texts from Wikipedia. Specifically, there are five categories in this framework, including: 1) ecology; 2) material culture; 3) social culture; 4) organizations, customs, ideas; 5) gestures and habits. Following a primary investigation of Wikipedia texts, we find that each entity-centered Wikipedia page is labeled by a variety of Wikipedia categories. To save the efforts of matching each entity on Wikipedia with each CSI category, we turn to map CSI categories with Wikipedia categories  Asthana and Halfaker (2018). To further guarantee the mapping quality, we manually create a mapping table for building connections on the categories from two resources. The resulting mapping table and category classification tool are described in Appendix 3.
2.2 Cultural Parallel Text Collection
To construct a cultural parallel text corpus (e.g., for En-Zh), we collect public text articles from Wikipedia’s translation tool that cover a wide range of cultural topics, and conduct sentence alignment to get a parallel corpus from that (The data collection details are in Appendix §C). To expand the language coverage in our corpora, we also reuse open-source parallel datasets from OPUS Tiedemann (2016). These include Wikipediav1.0 for English-French and English-Spanish, as well as Samanantarv0.2 for English-Hindi, English-Tamil, and English-Telugu. To identify cultural-specific sentences, we perform entity-linking  Ringgaard et al. (2017) to identify Wikipedia entities on the source texts, and use Wikiproject classification tool Asthana and Halfaker (2018) to classify these entities into cultural categories which are further mapped to our CSI categories using the cultural taxonomy (§2.1). Finally, we only keep the texts that contain items belonging to CSI categories.
2.3 Cultural Knowledge Augmentation
Existing MT studies Arthur et al. (2016); Hu et al. (2022) have been using external knowledge sources (e.g., Wikidata) to improve named entity translations. To enable future adaptations of these studies on our collected corpus, we parse Wikidata to extract the metadata of CSIs, which include their cultural labels, descriptions, and aliases in multiple languages. Furthermore, to identify items which are culture-specific, we collect information on the country of origin for each items and remove sentences containing items that do not have an associated origin country. This meticulous approach enabled us to enrich our dataset with supplementary information that can be utilized to evaluate the performance of machine translation models when handling culturally specific content.
2.4 Data Characteristic Summary
Pair Sent. CSIs Counts CSIs Types CSI Translations
En-Zh 778 794 601 730
En-Fr 2,073 2,213 2,213 1,130
En-Es 1,581 1,652 1,652 817
En-Hi 1,086 1,127 1,127 168
En-Ta 677 695 695 118
En-Te 754 695 695 66
Table 1: Dataset Statistics on Six Language Pairs.
Table 1 shows the statistics of our parallel corpora for the evaluation of MT systems on six language pairs. Particularly, for each language pair, we count the total number of detected CSIs by CSIs Counts and the number of unique CSIs by CSIs Types. It’s noted that not all the CSIs have translations on Wikidata, so we the number of CSIs containing the translations in WikiData by CSI Translations. Considering that many CSIs only exist within a specific culture group, which can’t be located in the parallel corpus, CSIs which don’t have a translation in other languages should take a higher proportion in real-world corpus than in our dataset.
3 Cultural Awareness Evaluation
Existing evaluation methods for terminology machine translation primarily focused on assessing the adequacy and fluency of the translated terms Anastasopoulos et al. (2021). However, CSI usually has more than one translation using various strategies, leading to challenges of fine-grained evaluation.
CSI-Match:
To evaluate the accuracy of translations, we introduce the CSI-Match metric as a modification to the traditional Exact Match (EM) evaluation metric. CSI-Match measures the accuracy of term translation using a more nuanced, fuzzy matching approach. CSI-Match is the maximal partial similarity ratio (PSR) between the reference CSI translations $\{t_{1},t_{2},\ldots,t_{n}\}$ and the system output sentence $S$. CSI-Match is calculated by Eq. (1), resulting in a value from 00 to $100$. A higher value indicates a stronger similarity of the predicted CSIs over a set of translation references.
[…]
PSR measures the maximum similarity between string $t$ and any substring in $S$.
$\text{PSR}(t,S)=\max_{s\in P}(1-d(t,s))\times 100$ (2)
$P=\{S_{i:j}\mid 0\leq i\leq j<|S|\}$ (3)
where $S_{i:j}$ is a substring of $S$ from word index $i$ to $j$, $d(,)$ is the normalized Levenshtein distance Levenshtein et al. (1966) between two strings.
Understandability in Human Evaluation:
In the context of evaluating CSI translation, we believe that, in addition to adequacy and fluency, the translated terms should also be comprehensible to speakers of the target language, considering that certain terms may not exist in the target language group. So, we design a human evaluation targeting at understandability which we define as the degree to which native target-language speakers can understand and explain the CSI translations. Specifically in the human evaluation, we ask the native target-language speaker to rank the understandability of CSI generated by various MT systems and evaluate the quality of CSI translation accordingly. To enhance our metric’s practical application, we leverage GPT-4 to compare the understandability between the target and MT systems, which has been shown as an effective way to evaluate the generation performances Rafailov et al. (2023). We calculate the understandability of each MT system as the win rate of the comparison. The evaluation prompt is detailed in Appendix §F.
Translation Strategies:
Furthermore, we categorize the translation strategies of CSI based on prior translation theories Newmark (1988); Persson (2015). These theories define different categorizations of strategies to improve the understandability of CSIs while keeping the cultural specialty. We select 4 strategies that are common to our dataset which are explained in Table 2. We then conduct a human evaluation to annotate the translation strategies employed by different MT models, and compare them with the strategies used for generating the reference translations.
Strategy Explanation
Transliteration Phonetically translates source CSIs
Literal translation Directly translates word-by-word
Description Integrates CSI descriptions of the CSIs
into the translation
Substitution Replace source CSIs by a semantically
equivalent item in the target language.
Table 2: Translation Strategies
4 Cultural Knowledge Prompting
In this section, we explore various prompting strategies to introduce cultural knowledge for LLM MT. We elucidate our strategies for generating in-context examples from external knowledge, which involve employing CSI translation pairs and CSI explanations. Additionally, we delve into several prompting strategies eliciting from LLMs’ internal knowledge. Table 3 shows examples of various prompting strategies.
Strategy Prompt
Basic Instruction (BI) Translate the following English text to Chinese
CSI Translation (CT) The Chinese translation of culture entities in the sentence is as following:
cannoli:里考塔芝士卷(Ricotta cheese rolls), 奶油甜馅煎饼卷 (Sweet Cream pancake rolls)
Translate the following English text to Chinese
CSI Explanation (CE) The explanation of culture entities in the sentence is as following:
Cannoli are Italian pastries consisting of tube-shaped shells of fried pastry dough …
Translate the following English text to Chinese
Self-Explanation (SE) User: Please explain cannoli in [Source Sentence]
LLM: [Explanation]
User: According to your explanations to cannoli, only translate the following English text to Chinese
Self-Ranking (SR) User: Translate the following English text to Chinese: [Source Sentence].
Please give [Generated Number] most likely translations, and ensure ”cannoli” in each result to correspond to different translations
LLM: [Translations]
User: Please select the best translation result of entities: cannoli. The translation of entitis in the sentence
should be the closest to its explanation, and is easy to be understood by Chinese readers.
Source They are also commonly available at Italian-American bakeries in the United States,
alongside other Italian pastries like cannoli and sfogliatelle.
Knowledge Translations:cannoli:里考塔芝士卷(Ricotta cheese rolls), 奶油甜馅煎饼卷 (Sweet Cream pancake rolls)
Explanation: Cannoli are Italian pastries consisting of tube-shaped shells of fried pastry dough …
Table 3: Prompting strategy examples (Top) and a source with cultural knowledge for En-Zh translation (Bottom)
4.1 External CSI Translation (CT)
We examine an external CSI dictionary for improving LLM-based MT. In particular, bilingual translation dictionaries play a vital role in the workflow of human translators Ghazvininejad et al. (2023). Here, we assess the impact of incorporating a CSI dictionary within the prompts. Specifically, we incorporate CSIs along with their corresponding translations prior to a basic translation instruction.
4.2 External CSI Explanation (CE)
CSIs may not have a direct equivalent in the target language’s culture when the concepts are not commonly used by the target language speakers. Therefore, it becomes necessary to translate based on the explanation of the CSI to assist the target audience in better understanding the content. To assess the impact of explanations, we include the CSI description obtained from Wikipedia in the prompt before the basic translation instructions. This allows us to study whether additional explanations of CSIs can enhance the MT performance.
4.3 Self-Explanation (SE)
We also examine LLMs’ internal knowledge for explaining the meaning of CSIs. Notably, chain-of-Thought (CoT) prompting has shown effective in eliciting LLMs’ internal knowledge Wei et al. (2022); Kojima et al. (2022). Inspired by this, we treat the explanation of CSIs in a source sentence as intermediate reasoning steps before translating the whole sentence. We then design the explanation prompting strategy in two steps for machine translation. In the first step, we prompt the LLM to explain the meaning of all CSIs in the source sentence. In the second step, we ask the LLM to translate the whole sentence by combining the LLM’s explanation with another prompt instruction.
4.4 Self-Ranking (SR)
Finally, LLMs can be prompted to sample different translation outputs. Even for multiple semantically equivalent prompts, the word choice in the prompts can have a significant impact on the performance of LLM models in translation tasks. To achieve more consistent and reliable results, we examine a self-ranking method Wu et al. (2023) to instruct LLMs to rank their translation outputs. Specifically, we prompt the model to first generate a fixed number of potential translations.
5 Experimental Settings
Methods in Comparison:
To fully evaluate the efficacy of LLM translations for cultural nuances, we compare the different prompting strategies (§4) on a tuning-free LLMs, as well as an open-sourced NMT model and a commercial MT system.
  • Prompting LLMs: We examine the different prompting strategies on OpenAI’s ChatGPT(gpt-3.5-turbo-1106) which is fine-tuned from GPT-3.5 with instructions. We also include LLaMA2-7B for comparison.
  • NLLB: We use the NLLB 1.3B Team et al. (2022) model, which is a state-of-the-art multilingual MT model.
  • Commercial MT: We use the Google Translate engine in our comparison.
[…]
6 Results and Analysis
Our analysis includes three parts: 1) automatic evaluations between LLM-based translation and popular NMT systems (§6.1) on six language pair translations; 2) assessing LLM prompting strategies using ChatGPT and LLaMA for En-to-Zh translation (§6.2); and 3) fine-grained human evaluation on a subset of En-to-Zh pairs (§6.3).
6.1 Overall Automatic Evaluation
(a) Traditional MT Systems
(b) LLM-based MT Systems
Figure 3: CSI-Match on Different Language Pair Translations: The coordinate scale range is from 40 to 100.
Model Overall-U NT-U Overall-U NT-U
EN-ZH ZH-EN
NLLB 22.4 17.7 6.38 0
NLLB-A 23.1 - 11.1 -
NLLB-R 22.9 - 17.3 -
LLaMA2 16.3 22.2 18.1 11.1
LLaMA2-A 20.2 - 29.3 -
LLaMA2-R 19.6 - 26.8 -
ChatGPT 55.8 61.9 54.4 72.2
Google 56.6 58.7 39.9 22.2
Table 4: Win Rates (%) of Understandability evaluated by GPT-4 for En-Zh Translation Pair: Overall-U means understandability on the entire dataset, and NT-U means non-translation CSIs’ understandability
Replacement and Append Work for NLLB and LLM-MT in CSI-Match.
Figure 3 shows the CSI-Match scores (§3) for eight methods across cultural-related parallel corpora in six language pairs, which are collected by our pipeline (§2). Figure 2(b) shows the CSI-Match scores of traditional MT systems including vanilla NLLB, NLLB-A, NLLB-R, and Google Translate, whereas Figure 2(a) shows scores of LLM-based MT systems including LLaMA2, LLaMA2-A, LLaMA2-R and ChatGPT. We find that straightforward strategies using dictionaries of CSIs, such as Replace (the green solid line) and Append (the orange solid line), are effective on metrics that rely on string-matching, such as CSI-match, as well as other semantic matching metrics detailed in Appendix §G. However, the appending strategy significantly benefits LLaMA more than NLLB. This suggests that LLaMA’s capacity of in-context learning and instruction-following enables the flexible integration of cultural knowledge at test time, a capability not presented in traditional NMT systems like NLLB. Additionally, we measure the CSI-Match of ChatGPT using basic instructions (BI), and Google Translate (the red solid line). We find that Google Translate’s performance across different languages is generally more consistent compared to ChatGPT. Specifically, Google Translate performs better in non-Latin languages such as Chinese, Hindi, Tamil, and Telugu, when compared to ChatGPT.
Understandability Gap on Non-translation CSIs
Table 4 presents the understandability win rates assessed by GPT-4 for various MT systems compared to the reference in English-Chinese translation. Appendix §F has the detailed evaluation prompt. The evaluation metrics include both the overall win rates across the entire dataset (Overall-U) and the win rates of samples containing complex source items (CSIs) with no translations (NT-U). The results indicate that traditional terminology translation methods, such as Appending and Replacing, effectively improve the overall understandability across the entire dataset. However, they still encounter challenges in improving the comprehensibility of translations that contain CSIs without a well-established reference. Notably, ChatGPT performs better than any other MT systems, which potentially suggests that the instruction-tuning of LLMs beyond the translation task may be beneficial for the model to generate explanation-based translations for these non-translation CSIs.
6.2 Prompting Strategy Evaluation
Shot Method CSI-Match Overall-U NT-U
Zero BI 67.2 55.8 61.9
CT 84.0 59.4 -
CE 66.6 55.5 58.7
SE 65.1 55.7 58.7
SR 69.2 58.5 47.6
Two CE 67.1 57.7 73.0
SE 67.7 58.7 58.7
SR 68.2 56.6 57.1
Table 5: Evaluation of different strategies by ChatGPT on English-Chinese translations.
As LLM-based MTs open up the opportunity to incorporate free-form external knowledge for improving the understandability of non-translation CSIs, we explore various prompting strategies by incorporating additional cultural information alongside dictionaries to improve translation. We compare different prompting strategies on ChatGPT, focusing on English-to-Chinese translation. We examine both zero-shot and 2-shot approaches to obtain results from ChatGPT. Especially for complex prompting strategies including CE, SE, and SR, we use 2-shot examples that are tailored for culturally-aware translation. Table 5 shows the results of different prompting methods.
Enternal Knowledge Prompting
Comparing the strategy of using external knowledge in prompts (i.e., CT and CE), we observe that LLMs can effectively leverage both direct translations and indirect descriptions. Specifically, CT greatly enhances the CSI-Match score for CSI translation. While 2-shot CE demonstrates limited improvements on CSI-Match, it improves the overall understandability of the entire dataset. More importantly, 2-shot CE significantly improves the non-translation CSIs’ understandability to 73.0. This suggests that CSI explanations offer significant assistance in translating CSIs, especially for the ones without well-known translations. In Table 6, we quantitatively show the output examples of “Polenta,” an Italian corn porridge. Interestingly, its Chinese translation on Wikidata is merely a transliteration. Under the CT strategy, ChatGPT directly copies this transliteration into the output, which may be considered correct yet not comprehensible for native speakers of the target language. In contrast, ChatGPT using CE integrates an explanation into the translation, describing the CSI as “corn porridge” in Chinese. This translation is much easier for readers to understand the nature of Polenta.
Internal Knowledge Prompting
Next, we compare prompting strategies to elicit LLMs’ internal knowledge (i.e., SE, SR). We find that 2-shot SE improves translation performance on CSI-Match and Overall-U but doesn’t improve the NT-U, which suggests that ChatGPT itself still lacks sufficient knowledge of non-translation CSIs which are typically very unique to the target-language culture. On the other hand, SR proves beneficial on ChatGPT for CSI-Match but does not improve the understandability of CSI translations, indicating that LLMs inherently possess internal knowledge of low-frequency CSIs but still can not capture cultural nuances through the SR prompting strategy examined in this study. In Table 6, 2-shot SR translates “Polenta” as “cornmeal,” similar to the baseline (BI). However, SE successfully translates it into “corn porridge.” This suggests that ChatGPT possesses background knowledge of Polenta but struggles to spontaneously elicit it to generate and select the most understandable translations.
Strategy Outputs
BI 玉米面 (cornmeal)来自意大利北部,遍布阿根廷。
CT 波伦塔 (transliteration) 起源于意大利北部,在阿根廷非常常见。
CE (2-shot) 玉米粥 (corn porridge) 源自意大利北部,是阿根廷非常常见的食品。
SE (2-shot) 玉米粥 (corn porridge) 起源于意大利北部,是在阿根廷非常常见的食物。
SR (2-shot) 玉米面 (cornmeal) 来自意大利北部,在阿根廷非常常见。
Source Polenta comes from Northern Italy and is very common throughout
Argentina
Reference 玉米粥 (corn porridge)来自意大利北部,在整个阿根廷很常见
Knowledge Translation: 波伦塔 (transliteration)
Explanation: Polenta is a dish of boiled cornmeal that was historically
made from other grains. The dish comes from Italy. It may be served
served as a hot porridge.
Table 6: Output Examples of Prompting Strategy (Top) and a source-reference sentence pair with cultural knowledge for En-Zh translation (Bottom)
6.3 Human Evaluation
GPT-4 and Human Evaluation Are Consistent.
Table 7 shows the Overall-U evaluated by both GPT-4 and human annotators. Specifically, we randomly select 200 samples from the English-to-Chinese translation dataset and have human annotators evaluate the understandability of output translations by the four different methods. The results indicate a high level of consistency between human annotations and GPT-4’s evaluations
Method GPT-4 Overall-U Human Overall-U
NLLB 19.0 16.0
LLaMA2 13.0 14.0
ChatGPT 52.5 51.0
Google 54.0 55.5
Table 7: GPT-4 v.s. Human in Overall-U evaluation
Translation Strategies
To explore potential factors benefiting understandability, we let a human annotator examine the models’ translation strategies. In addition to the four strategies (mentioned in §3; Table 2), we propose three extra considerations 1) Copy that directly copies the source language of CSI into the target language; 2) Wrong that indicates entirely incorrect translations; and 3) Other that employs other strategies in translation. Figure 4 shows the ratio of each strategy in four MT systems. We find that models with higher understandability (i.e., ChatGPT and Google translate) use description and substitution at a significantly higher rate, indicating that these two strategies help improve the understanding of CSI for target-language speakers. Notably, LLaMA2 incorporates a higher frequency of substitution and description methods compared to traditional NLLB. However, this increased diversity in translation output comes at the cost of reduced stability in the outputs. As a result, LLaMA2 tends to yield more inaccurate translations, whereas NLLB relies more on Literal Translation and Transliteration to translate CSIs.
To compare the consistency between automatic and human evaluation, we also calculate the correlation between automatic evaluation metrics and human evaluation, where we consider the translation correct in human evaluation if the translation is categorized into all the translation strategies except Copy and Wrong. The Pearson’s correlation between the human evaluation and BLEU, BLEURT and COMET are 88.8, 90.5, and 89.5 respectively, whereas CSI-Match has the highest correlation score at 94.7 with human evaluation, suggesting that CSI-Match can be an efficient evaluation metric for CSIs with a reference translation.
Figure 4: Percentage of Translation Strategies
7 Related Works
Cultural-aware Machine Translation:
As languages and cultures are highly intertwined, there is a growing desire towards empowering cultural awareness of machine translation systems Nitta (1986); Ostler (1999); Hershcovich et al. (2022). However, as cultural nuances are subtle, collecting culturally sensitive data Akinade et al. (2023) remains costly and time-consuming. Besides, it is also challenging to perform a human-centered evaluation of the cultural nuances Liebling et al. (2022). Existing studies have proposed strategies to evaluate cultural awareness of traditional MT systems by grounding images Khani et al. (2021), adapting entities Peskov et al. (2021) or targeting at dilates Riley et al. (2022). Different from evaluating traditional MT systems, we focus on evaluating the cultural awareness of LLM-based translation.
MT with Terminology Dictionaries
Previous studies have proposed to integrate terminology dictionaries into NMT models Dinu et al. (2019) and LLM-based MT systems Ghazvininejad et al. (2023), proving effective for terminology translation. A major thrust of methods either modify the model architectures to integrate the dictionary or feed the translation dictionary as parts of inputs to NMT models. In this study, we focus on evaluating LLM-based MT methods with conventional NMT methods without modifying the underlining models, as the parameters of LLMs (e.g., ChatGPT) may not be accessible. Moreover, compared to traditional terminology translation from popular domains (e.g., finance, medicine), translating cultural-specific items carries its unique challenge. CSIs are unique to specific cultural groups, leading to difficulty in understanding the cultural nuance of CSIs for individuals from other cultures. Thus, beyond accuracy-focused metrics, assessing CSI understandability is also crucial and underexplored in the literature.
LLM-based MT:
Large language models, such as GPT-3 Brown et al. (2020), have proven effective in machine translation for various high-resource languages Hendy et al. (2023); Jiao et al. (2023). Particularly, a few recent studies have investigated the performance of LLM-based MT, including formality control of translation outputs Garcia and Firat (2022), in-context translation ability during pre-training Shin et al. (2022), and multilingual translation Scao et al. (2022). However, the exploration of LLM-based MT on culturally sensitive texts is still missing.
External Knowledge for MT:
There have been multiple threads of research efforts on integrating external knowledge such as bilingual translation lexicons for neural machine translation systems, including probability interpolation of lexicons Arthur et al. (2016); Khandelwal et al. (2021), data augmentation by back-translations Hu et al. (2019), decoding with a phrase memory Wang et al. (2017), and pre-training with an entity-based denoising objective Hu et al. (2022). Despite the effectiveness, these methods require further fine-tuning of the original MT systems, while we focus on tuning-free methods for integrating external knowledge for LLM-based MT in this study.
8 Conclusion
In this paper, we introduce a novel data curation pipeline aimed at constructing a culturally sensitive parallel corpus to assess the cultural awareness of MT systems. Additionally, we propose a reference-free metric to evaluate the understandability of translations for cultural-specific content by GPT-4. We also devise simple yet effective prompting strategies to enhance the understandability of LLM-based translations. Despite the effectiveness, several challenging questions remain open. First, it is non-trivial to incorporate cultural-specific information beyond a single entity such as discourse information. Besides, our prompting strategies leverage mostly external cultural knowledge in the form of texts. How to leverage multimodal knowledge from images and structured knowledge graphs to resolve cultural ambiguity deserves further investigation.
Limitations
Our work provides one step further toward understanding and evaluating the cultural awareness of LLM-based machine translation. We provide a culturally sensitive parallel corpus with rich annotations on cultural-specific items. Thereby, prompting strategies in our study, due to the cost ChatGPT API, is limited to the language pair (i.e., English-Chinese). To manually verify the quality of the data, we control the data size in a small one. We will provide our code repository to facilitate the future adaptation of our pipeline for more language pairs in diverse language families. Besides, as we focus on the evaluation of cultural-specific items in this study, the evaluation of cultural awareness beyond a single entity deserves further investigation. In addition to cultural-specific items, there are many other types of cultural errors that exist in the translation process, such as linguistic style and slang Hershcovich et al. (2022). Our work aims to mitigate cultural errors by starting from CSI, in order to promote advancements in cultural-aware machine translation datasets, models, and evaluation methods. This is of great importance for enabling machine translation to play a larger role in cross-cultural contexts. Further, we only try several promoting strategies in our study, which is because our work focuses on benchmarking the cultural awareness of current LLM-based MT systems, in the future, we’ll test other methods such as instruction tuning to improve the performance.
Ethical Considerations
Although our study designs a suite of simple but effective prompting strategies to enhance the cultural awareness of LLM-based machine translation, we still observe the weakness of LLM-based translation on cultural concepts in certain regions (e.g., Asia) and hallucinations on low-frequency entities. Potential usage of these LLM translation outputs may still result in misinformation spread. Before deploying our methods to create reliable content such as creating translations of Wikipedia articles, practitioners should ensure another round of human post-editing. In the future, we’ll make our dataset public to assist further research on cultural-awareness of MT systems. During the annotation process, the annotators (native speakers of the target languages) consist of the authors of this article, who know the goals of the study.
Appendix A Data Examples
In Table 8, we present a data example from the English-Chinese corpus. Each data point consists of a pair of sentences. We meticulously annotate all cultural-specific items (CSI) within the sentences. For each cultural-specific item, we provide information including its category, country of origin, translations in the target language, descriptions in both the source and target languages, and an explanation. To illustrate the challenges that cultural-specific items pose for current Machine Translation (MT) systems, we provide translations from both Google Translate and ChatGPT for this example. It is noted that both Google and ChatGPT erroneously rendered the Chinese translation of ”Wiener Schnitzel” as ”pork chops” instead of the correct translation, which is ”steak.” This misinterpretation not only misleads Chinese readers but also introduces confusion to the entire sentence.
Aspect Content
Source (EN) The Shanghai-style Fried Pork Chop is
a modification from Wiener Schnitzel
the national dish of Austria, and a fried pork
chop is more a street food than a beef steak.
Target (ZH) 上海炸猪排的做法改良自奥地利国菜
维也纳炸牛排 (Wiener fried steak),而炸猪排
与牛排不同,它显得更加市井。
Cultural-Specific Item Wiener Schnitzel
Category Culture.Food and drink
Country of Origin Austria
Translation (ZH) 维也纳炸牛排 (Wiener fried steak)
Description (EN) breaded veal schnitzel
Description (ZH) 面包屑小牛肉炸肉排
Explanation The entity, sometimes spelled Wienerschnitzel,
is a type of schnitzel made of a thin,
breaded, pan-fried veal cutlet. It is one of
the best known specialities of Viennese cuisine,
and one of the national dishes of Austria.
NLLB 上海风格的炸猪肉切片是从奥地利
国家菜的维也纳施尼切尔(transliteration)改制而成,
LLaMA2 上海炒猪排是一种来自奥地利
牛肉炒肉块 (Beef stir-fried cubes)的改良型,
而炒猪排更像是一道街头小吃而非牛肉炒肉块.
Google Translate 海派炸猪排是奥地利国菜
维也纳炸猪排 (Wiener fried pork chops)
的改良版,炸猪排与其说是牛排,
不如说是街头小吃。
ChatGPT 上海式炸猪排是从奥地利的国菜
维也纳炸猪排 (Wiener fried pork chops)
改编而来,而炸猪排比牛排,
更像是街边食物。
Table 8: A Data Example of the English-Chinese Corpus: In parentheses, we explain what the Chinese translation means.
Appendix B CSI vs. Wikiproject Mapping Table
The mapping table between CSI definition (5 categories in all) and Wikiproject categories [^2] https://en.wikipedia.org/wiki/Wikipedia:WikiProject_Categories (18 categories) are shown in Table 9. Additionally, we provide illustrative examples for each category to clarify the respective meanings. The tool we used for Wikiproject category classification is drafttopic [^3] https://github.com/wikimedia/drafttopic.
CSI Category Wikiproject Category CSI Example
Material Culture Culture.Food and drink Cotoletta (Italy)
Culture.Visual arts.Architecture the Summer Palace (China)
History and Society.Transportation Kan-Etsu Expressway (Japan)
Social Culture Culture.Sports RKC Waalwijk (Netherlands)
Culture.Media.Entertainment Far Rockaway (USA)
Organisations, History and Society.Politics and government Europe Ecology – The Greens (France)
Customs and Ideas Culture.Philosophy and Religion Fuller Theological Seminary (USA)
Culture.Literature Der Spiegel (German)
Culture.Visual arts.Visual arts* The Headless Horseman Pursuing Ichabod Crane (USA)
Culture.Visual arts.Fashion Bottega Veneta (Italy)
Culture.Visual arts.Comics and Anime Dragon Ball (Japan)
Culture.Performing arts Just Dance (USA)
Culture.Media.Music Trident Studios (UK)
Culture.Media.Films A Few Good Men (USA)
Culture.Media.Books Moby-Dick (USA)
History and Society.History Tusculum (Italy)
Ecology STEM.Biology Kapok (Netherlands)
Geography.Regions.* Qualicum Beach (Canada)
Gestures and Habits -
Table 9: CSI vs. Wikiproject Mapping Table
Appendix C Wikipedia Parallel Corpus Collection
To collect English-Chinese parallel corpus from Wikipedia, we use the bilingual Wikipedia articles translated through Wikipedia’s content translate tool [^4] https://en.wikipedia.org/wiki/Wikipedia:Content_translation_tool. This tool allows confirmed editors to translate Wikipedia articles from a source language to a target language with a machine translation system. By tracking their editing logs, we obtain the text triples consisting of the original text in a source language, the machine-translated text, and the human post-edited text in a target language. We then use a sentence alignment tool bleu-align [^5] https://github.com/rsennrich/Bleualign Sennrich and Volk (2010) to obtain a sentence-level parallel corpus. Then we use open-source data from OPUS, which includes Wikipediav1.0  [^6] https://opus.nlpl.eu/Wikipedia-v1.0.php for English-French and English-Spanish, as well as Samanantarv0.2  [^7] https://opus.nlpl.eu/Samanantar-v0.2.php for English-Hindi, English-Tamil, and English-Telugu.
Appendix D Data Characteristics
Figure 5: Data characteristics on regions (Outside) and categories (Inside) for the English-Chinese corpus
Culture is intricately linked with specific regions, and its manifestations can exhibit substantial variations across diverse regions and categories. Therefore, our dataset encompasses cultural-specific elements sourced from a wide array of regions and categories, as shown in Figure 5, which shows the distribution of the English-Chinese corpus. This inclusive approach allows us to comprehensively evaluate the performance of machine translation models across a broad spectrum of cultural contexts.
Appendix E Experiment Settings
The experiment settings of different models included in our paper are as follows:
Appendix F Evaluation Prompts of GPT-4
It’s shown that GPT-4 could be an effective tool for evaluating the generation task’s quality in DPO, we try the same prompt manner with the edition towards understandability of CSIs, the prompt is as follows:
Assuming you’re a Chinese native speaker, which of the following translations does a better job of translating the CSI word “Goubuli” into understandable Chinese words?
¡Source¿
Translation A: ¡A¿
Translation B: ¡B¿
FIRST provide a one-sentence comparison of the two translations, explaining which you prefer and why. SECOND, on a new line, state only ”A” or ”B” to indicate your choice. Your response should use the format:
Comparison: ¡one-sentence comparison and explanation¿
Preferred: ¡”A” or ”B”¿
Appendix G Overall Automatic Evaluation
Language Pair Method BLEU BLEURT COMET CSI-Match BLEU BLEURT COMET CSI-Match
English-Chinese EN-ZH ZH-EN
NLLB 23.2 0.558 77.0 53.1 27.0 0.594 76.5 64.9
NLLB-A 17.3 0.447 62.8 58.3 21.9 0.531 69.4 65.5
NLLB-R 23.9 0.555 76.8 78.7 25.3 0.591 75.4 79.8
LLaMA2 17.1 0.529 75.8 45.0 26.1 0.595 77.9 70.9
LLaMA2-A 18.3 0.518 74.4 80.2 29.1 0.629 79.0 85.6
LLaMA2-R 19.2 0.547 76.8 78.1 27.7 0.618 78.9 80.6
ChatGPT 29.3 0.642 82.9 67.2 32.5 0.668 80.9 77.7
Google 38.3 0.679 84.2 71.9 41.1 0.697 82.2 79.5
English-French EN-FR FR-EN
NLLB 37.4 0.585 78.3 77.7 36.3 0.634 77.8 88.9
NLLB-A 37.4 0.582 78.0 77.4 35.4 0.628 77.2 88.3
NLLB-R 37.0 0.577 77.8 92.6 36.6 0.635 77.7 92.1
LLaMA2 34.4 0.558 76.2 77.9 27.6 0.462 66.0 72.5
LLaMA2-A 35.0 0.568 67.8 82.7 34.7 0.633 77.7 92.8
LLaMA2-R 34.7 0.550 75.6 89.8 30.2 0.620 76.8 93.3
ChatGPT 36.2 0.594 78.9 78.7 37.6 0.629 77.8 89.8
Google 31.1 0.573 77.5 77.9 36.1 0.632 77.2 88.9
English-Spanish EN-ES ES-EN
NLLB 48.8 0.707 83.4 78.4 50.7 0.718 83.5 90.3
NLLB-A 48.3 0.705 83.3 78.7 49.9 0.713 83.0 90.8
NLLB-R 47.9 0.696 82.9 94.0 50.9 0.717 83.4 95.2
LLaMA2 44.6 0.679 81.6 76.9 45.3 0.704 82.6 89.9
LLaMA2-A 44.5 0.674 81.2 82.0 46.6 0.706 82.7 93.3
LLaMA2-R 44.5 0.675 81.3 93.0 44.2 0.702 82.3 94.7
ChatGPT 47.9 0.711 83.6 79.3 50.7 0.712 83.4 92.3
Google 42.9 0.704 82.0 79.1 49.8 0.722 83.0 90.9
English-Hindi EN-HI HI-EN
NLLB 32.8 0.637 0.747 73.9 38.4 0.683 83.5 93.7
NLLB-A 32.1 0.630 0.734 78.7 38.4 0.676 82.8 90.9
NLLB-R 32.9 0.639 0.748 83.6 38.4 0.684 83.5 98.3
LLaMA2 7.3 0.438 0.493 60.6 12.6 0.441 63.8 67.5
LLaMA2-A 7.0 0.439 0.493 81.9 18.9 0.546 74.3 79.1
LLaMA2-R 8.2 0.444 0.502 80.0 14.8 0.480 67.7 67.7
ChatGPT 24.1 0.592 0.701 72.2 32.0 0.649 81.4 93.4
Google 33.8 0.651 0.753 77.2 39.9 0.690 83.0 94.5
English-Tamil EN-TA TA-EN
NLLB 25.8 0.706 83.0 64.4 31.5 0.645 80.4 92.1
NLLB-A 25.5 0.698 82.3 70.5 31.0 0.639 79.8 94.9
NLLB-R 25.8 0.707 82.9 81.6 31.5 0.645 80.4 97.9
LLaMA2 1.7 0.309 39.6 44.0 4.3 0.331 51.9 54.4
LLaMA2-A 1.2 0.341 39.6 88.9 9.1 0.417 61.5 87.0
LLaMA2-R 1.4 0.321 40.0 73.9 5.1 0.305 53.7 68.7
ChatGPT 10.4 0.496 62.5 62.6 16.9 0.539 72.2 87.4
Google 26.9 0.712 82.7 68.7 31.4 0.649 80.4 91.9
English-Telugu EN-TE TE-EN
NLLB 31.4 0.628 81.1 80.6 34.7 0.643 81.2 87.3
NLLB-A 29.2 0.624 80.4 81.9 34.5 0.635 80.5 89.8
NLLB-R 31.4 0.628 81.1 89.8 34.7 0.643 81.2 94.7
LLaMA2 3.2 0.207 41.0 52.0 0.9 0.190 41.6 33.2
LLaMA2-A 4.0 0.244 42.3 88.8 5.8 0.356 56.7 78.0
LLaMA2-R 4.0 0.237 42.0 77.0 2.3 0.269 48.8 44.1
ChatGPT 16.8 0.484 67.3 69.3 23.3 0.567 74.8 78.8
Google 32.7 0.635 81.0 81.6 34.9 0.653 81.2 89.5
Table 10: Automatic evaluation of all MT methods on six language pairs from both translation directions.
Automatic Evaluation Metrics:
We first evaluate the translation outputs using traditional automatic metrics such as BLEU Papineni et al. (2002), BLEURT Sellam et al. (2020), and COMET Rei et al. (2020). To be consistent with the evaluation method of NLLB, we calculate spBLEU Goyal et al. (2022) for BLEU scores. In addition to traditional machine translation evaluation metrics, we also use CSI-Match to evaluate the translation quality of CSIs (described in §3). Table 10 shows the comparison results of six MT methods across six language pairs in both directions.
As shown in Table 10, NLLB displays high performance in both directions of translation over six language pairs. Both CSI dictionary incorporation (NLLB-A) and term replacement strategies (NLLB-R) enhance the translation quality of CSI for most language pairs, without significantly compromising the overall sentence translation regarding other metrics. Notably, NLLB-R outperforms other MT systems on CSI-Match, even including LLM-based MT. Interestingly, LLaMA2-7B shows an obvious drop in both traditional evaluation metrics and CSI-Match scores when translating English to three Indian languages and vice versa. One possible explanation is because of the insufficient Indian data during the pre-training of LLaMA2. Both CSI-involving translation strategies are beneficial for LLaMA-based translation. In non-Latin languages (i.e., Chinese, Hindi, Tamil, and Telugu), LLaMA2-A tends to yield better performances, whereas LLaMA2-R performs better in Latin languages (i.e., French and Spanish), which potentially suggests that injecting cultural knowledge through code-switching similar Latin languages works better than distant languages for LLM-based models. Furthermore, we assess the translation performances of ChatGPT and Google Translate. Both MT systems exhibit commendable performance in CSI translation, with Google Translate demonstrating superior translation results. Notably, Google Translate showcases consistent translation abilities, particularly in handling relatively low-resource languages like Tamil and Telugu.
Appendix H Automatic Evaluation of Prompting Strategies on LLaMA2
while on LLaMA2, we employed a 2-shot prompting method, which is different from the previous 8-shot in-context learning (§6.1) due to the LLaMa2-7B input length limit. The understandability of LLaMA-2 is shown in Table 11. 2-shot prompting strategies can hardly improve the understandability of LLaMA2-7B. The complexity of prompts even causes unstable generation results which show worse performance than baseline.
Shot Method CSI-Match Overall-U NT-U
two BI 35.1 16.3 22.2
CT 70.3 16.0 -
CE 35.5 11.1 14.2
SE 35.1 11.8 15.8
SR 37.9 10.2 11.1
Table 11: Evaluation of different strategies by LLaMA2 on English-Chinese translations.
Appendix I Performance Across Regions
Figure 6: Avg.CSI-Match by Regions
Culture is often associated with a specific region, and its expressions can vary significantly across different regions and categories. To gain a deeper understanding of the influence of region on CSI translation, we categorized CSI into three groups: CSI originating from countries primarily using the source language, countries predominantly using the target language, and countries utilizing languages other than the source and target languages. On the six groups of English-to-XX translations, we calculated the average CSI-Match values of these three CSI groups respectively, and the results are shown in Figure 6.
Given that target CSIs must have the translation in the target language. Translating target CSIs is akin to back translation. However, when translating the source CSI or other CSIs, the translation may either not exist in the target language or exist with lower word frequency. Consequently, the model is expected to yield better results for the target CSI. Surprisingly, our analysis reveals that most models excel at translating the target CSI back into the target language in Latin languages (i.e., French and Spanish). Notably, Google Translate consistently achieves superior translations across all languages. ChatGPT demonstrate better translation performance in Chinese and Tamil, while LLaMA2 succeeds in Hindi and Tamil for target CSI translation. In contrast, traditional translation models NLLB struggle with all non-Latin languages, failing to outperform the source CSI translation. This suggests that LLMs may possess enhanced learning capabilities for translating culture-related content. However, it is important to note that the current translation performance is not consistently stable.
Appendix J Comparison of Translation Strategies
Figure 7: Comparison of Translation Strategies: The value in each grid represents the winning rate of the method on the x-axis in comparison to the method on the y-axis.
In order to further compare the impact of different translation strategies on understandablity, we analyzed the comparison between different translation strategies based on the ranking results of human evaluation. Specifically, we also rank the different translation strategies used by the MT systems according to the rank of the MT system’s understandablity given by humans, which is shown in Figure 7. It’s shon that the winning rate of descriptions for all methods surpasses 0.5, and the winning rate of substitutions, excluding descriptions, significantly exceeds 0.5. This implies that translations employing these two strategies are generally deemed more comprehensible by human annotators. Moreover, Literal Translation outperforms Transliteration, highlighting that transliteration may diminish the clarity of CSI in translation compared to a literal approach. Notably, the winning rate of copying for both Literal Translation and Transliteration hovers around 50%, indicating that these two methods may introduce confusion, and their readability underperforms directly copying the original word.